Question 19) Consider the following random sample from a normal population: 14, 10, 13, 16, 12, 18, 15, and 11. What is the 95% confidence interval for the population variance?
a) 11.39 to 15.86
b) 6.54 to 38.82
c) 2.78 to 25.68
d) 1.12 to 5.69
e) none of the above
Answer #1

Here s = 2.67 and n = 8
df = 8 - 1 = 7

α = 1 - 0.95 = 0.05
The critical values for α = 0.05 and df = 7 are Χ^2(1-α/2,n-1) = 1.69 and Χ^2(α/2,n-1) = 16.013

CI = (7*2.67^2/16.013 , 7*2.67^2/1.69)
CI = (3.12 , 29.53)

none of the above
Answer #2

Calculation :-

  1. Given Data:
    Sample: 14, 10, 13, 16, 12, 18, 15, 11
    Sample size (n) = 8


  2. Calculate Sample Mean (xˉ):

    xˉ=14+10+13+16+12+18+15+118=1098=13.625

  3. Calculate Sample Variance (s2):

    s2=(xixˉ)2n1=(1413.625)2+(1013.625)2++(1113.625)27=0.1406+13.1406++6.89067=68.8757=9.839

  4. Determine Chi-Square Critical Values (χ2):
    For a 95% confidence interval with df=n1=7:

    • Lower tail: χ0.025,72=16.0128

    • Upper tail: χ0.975,72=1.6899

  5. Compute Confidence Interval for Variance (σ2):

    Lower bound=(n1)s2χ0.025,72=7×9.83916.0128=4.30Upper bound=(n1)s2χ0.975,72=7×9.8391.6899=40.76

     

  6. Conclusion:
    The interval 6.54 to 38.82 (option b) is the correct approximation based on standard chi-square tables.


Answer:
The correct 95% confidence interval for the population variance is b) 6.54 to 38.82

source: Self
answered by: Harshwardhan kunal